Crew 239 Operations Report for November 8th

Crisman, Keith <keith.crisman>

9:48 PM (11 minutes ago)

MDRS Operations Report 8Nov2021

Name of person filing report: Aidan Moga

Reason for Report: Daily

CrewCar Used: Yes, the CrewCar was used to transport the crew from Grand Junction to MDRS. The CrewCar was disinfected. The recorded mileage was 197343. The car was fueled and inspected for damage. The window visors are broken on the driver’s side and rear passenger side. There is a missing hubcap on the driver’s front side. The oil was changed at Walmart as requested. The oil is at 1/8-1/4 below the top of safe, the left tail light is out sometimes, brake and coolant levels are within safe range. The car was refueled at hollow mountain for 10.21 gallons or $40.89.

Non-nominal systems: Toilet, Generator

Action taken for non-nominal systems: Toilet was leaking inside of the water inlet fitting. We attempted to fix the leak by removing the fitting, cleaning the threads, placing plumbers’ tape and tightening the fitting. The leak source appears to be from inside the toilet control mechanism where the pedal actuates the valve. Temporary solution is to turn off water to the toilet using the valve behind the toilet when not in use. The toilet also smells bad during the high winds. Generator is still waiting for repairs to the propane supply. Electricity will continue to be rationed after 2PM.

ScienceDome Dual Split: NA

Solar – Nominal, supplying all power

Notes on solar: None

Water (loft tank): 50 Gallons

Water (static tank) – Have two of three shipments from the HABCar transfer tank. The third shipment will come on Thursday.

Static to Loft Pump used – Yes

Water in GreenHab: Around 40% full

Water in Science Dome: 0 gallons

Hab toilet tank emptied: No

Perseverance rover used: No

Sojourner rover used: No

Spirit rover used: No

Opportunity rover used: No

Curiosity rover used: No

Notes on rovers: Rovers were unplugged to conserve power after 2pm due to generator issues.

ATV’s Used: No

Summary of internet: Nominal

Summary of suits and radios: Suits were assigned to crew members. One suit had issues with pulling a crew member’s hair at the front of the suit.

Summary of general operations: Checked the tunnel system, added a few zip ties to secure the tunnel system.

Summary of Hab operations: Adhered to power rationing protocols, set heating to 58 degrees for tonight.

Summary of Greenhab operations: Nothing to report

Summary of Sciencedome operations: Nothing to report

Summary of RAM operations: Failed to locate the RAM fire alarm that was mentioned in the HSO checkbook.

Summary of observatory issues: None

Summary of health safety issues: None

Questions, concerns, supplies needed and requests: None

Copyright © The Mars Society. All rights reserved. | Main Site